OpenStudy (twizttiez):

What is the solution to the system of equations represented by the two equations? y=2/3x y=-2/3x+4

OpenStudy (twizttiez):

@Write0The0Right0For0Wrong0Doing

OpenStudy (anonymous):

u can help me after these:) with english

OpenStudy (twizttiez):

Will do

OpenStudy (anonymous):

no pic for this one?

OpenStudy (twizttiez):

Nope

OpenStudy (anonymous):

ugh

OpenStudy (twizttiez):

Can you help?

OpenStudy (twizttiez):

@funfuse2

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

is it a multiple choice answer?

OpenStudy (twizttiez):

I could use your help

OpenStudy (anonymous):

is it

OpenStudy (princeharryyy):

y=2, x=1/3

OpenStudy (twizttiez):

No What is the solution to the system of equations represented by the two equations? y=2/3x y=-2/3x+4 (___),(___)

OpenStudy (princeharryyy):

(1/3,2)

OpenStudy (twizttiez):

Thank you i need help on a few more could you help?

OpenStudy (princeharryyy):

k

OpenStudy (twizttiez):

@princeharryyy i cant type in 1/3 on this

OpenStudy (princeharryyy):

i'll give d answers i hate typing actually

OpenStudy (princeharryyy):

on what

OpenStudy (princeharryyy):

try 3^(-1) they are equivalent

OpenStudy (princeharryyy):

@TwiztTiez

OpenStudy (princeharryyy):

n plz be quick i'm feeling sleepy its alredy 12:30 am here

OpenStudy (twizttiez):

1 moment i have to do some writing here is another ?

OpenStudy (princeharryyy):

k